OpenStudy (anonymous):

Kerri correctly calculated the distance between point S (-3428) and point T (286) on a number line. What value did Kerri find for the distance? A. -3714 B. -3142 C. 3142 D. 3714

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

easy peazy. If you start from -3428 then how do you think you'd get to 0? Go ahead and give it a try.

OpenStudy (anonymous):

You still there?

OpenStudy (anonymous):

Yep. I'm working on it!

OpenStudy (anonymous):

Alright take your time.

OpenStudy (anonymous):

If you are feeling overwhelmed by that number then try a number like -2. How would you get from -2 to 0?

OpenStudy (anonymous):

I don't have all day.

OpenStudy (anonymous):

meh, I have to leave but I'll just explain it for you. To get to 0 from -3428 all you have to do is add the same number to itself. So like if I had -2 and I added 2 to it I'd get 0. So If I say -3428 + 3428 = 0 and 0 + 286 = 286 then 3428 + 286 = the distance. I think you probably have basic math skills so I'll let you add the two numbers together.

OpenStudy (anonymous):

I got it thanks!

OpenStudy (anonymous):

You forgot to give me a medal...

OpenStudy (anonymous):

i gave you one @thedogkeeper